OpenStudy (anonymous):

Help me please? Simplify the following:

OpenStudy (anonymous):

\[\sqrt{108x4y9}\]

OpenStudy (anonymous):

I think you mean? \[\sqrt{108x^4y^9}\] find a perfect square factor in 108 first. You can split 108 to 36 and 3, and 36 is a square. So take the square root of 36 and put it outside the square root. \[6 * \sqrt{3x^4y^9}\]

OpenStudy (anonymous):

than, take the square root of even exponents out. Odd exponents, leave one inside. \[6 * x^2 * y^4 \sqrt{3y}\]

OpenStudy (anonymous):

so \[6x^2y^4\sqrt{3y}\]